OpenStudy (anonymous):

how the lim (ln(2+n)/n)/(1/n) = lim (n/2+n) * (n-(2+n)/n^2 ) / (-1/n^2)....[ lim is going to + infinity]

OpenStudy (anonymous):

find the value of c if \[\sum_{2}^{\infty}\](1+c)^ -n = 2

OpenStudy (myname):

\[\sum_{2}^{\infty}1/(1+c)^n =\sum_{0}^{\infty}1/(1+c)^n - \sum_{0}^{2}1/(1+c)^n---(1)\] \[\sum_{0}^{\infty}1/(1+c)^n = 1/1-(1/(1+c))\] [BY USING THE FORMULA TO FIND THE SUM OF THE GEOMETRIC SERIES-->a/1-r ] Then we simplify, \[\sum_{0}^{\infty}1/(1+c)^n = (1+c)/c\] Next, \[\sum_{0}^{2}1/(1+c)^n=1+(1/1+c) = (2+c)/(1+c)\]Now we can plug in the values of these series in (1). \[2 = ((1+c)/c) - (2+c)/(1+c)\] Now, take the common denominator. \[2= ((1+c)^2-2c-c^2)/c(1+c)\] \[2c +2c^2= 1+2c+c^2-2c-c^2\] \[2c+2c^2-1=0\] Here , you can use the quadratic formula to solve for c.[c is a variable like x] By doing that you would get c =0.366 and c= -1.366 You can go head and plug the two c values in the series , and you will find which one gives you 2 and which one does not. In this case c is 0.366